Given that cscθ=sec0.75 find θ in radians rounded to 2 dec. pl.
Answer by lwsshak3(6505)

Given that cscθ=sec0.75 find θ in radians rounded to 2 dec. pl.
**
sec=1/cos
sec(.75)=1/cos(.75)=1.37
θ=arccsc(1.37)=arcsin(1/1.37)=.82 radians
ans:
θ=0.82 radians